A 74 Be atom (mass = 7.0169 amu) decays into a 73 Li atom (mass = 7.0160 amu) by electron capture. How much energy (in millions of electron volts, MeV) is produced by this reaction?
Answer #1

Sol:the electron capture reaction is shown below.

74Be + -1e ->73Li

Mass defect(∆m)=mass of Be+mass of e - mass of Li

=7.0169+0.00054858-7.0160 amu

=1.449×10-3amu

=1.449×10-3amu

E=∆m×931 MeV

=1.449×10^-3×931 MeV

=1.349MeV